# Motion Capture from Inertial and Vision Sensors

### v3

@@ I Introduction
Human motion capture is the process of recording human movement represented by a sequence of 3D positions and rotations of mesh or joints of the human body [26, 43, 44]. Industrial motion capture systems have been widely applied in movie and game production, sports analysis, medical diagnosis, etc. However, these systems usually consist of tens of synchronized cameras or a group of wearable sensors and specific signal receivers [1, 2, 56]. Despite their high accuracy for human motion capture, individual consumers can hardly afford the high cost and learn professional configurations. Therefore, this paper investigates accurate multi-modal human motion capture with consumer-affordable devices and easy-to-use operations for daily applications like eXtended Reality (XR), mobile video production, live video streaming, etc.
@@
Based on MINIONS, we propose a two-branch framework, named SparseNet, for motion capture by fusion of videos and IMUs. The core of SparseNet is a deliberately designed two-branch structure grounded in Bayesian probability theory and Matrix Fisher priors [13, 30]. For the sparse IMU signals, the Sparse IMUs branch takes inertial measurements as input to learn posture features, providing rotational probability distribution. To overcome the incompleteness and location drift of sparse IMUs, we design a visual branch to supplement the missed joints and correct the location drift, estimating 3D bone probability distribution and pose probability prior. This framework enables adaptive integration through a posterior fusion module that aligns distributions from both branches. Experiments reveal that the visual branch for monocular motion capture is often limited by jittering due to blurriness or occlusion, while the IMUs branch tends to suffer from global position drifts. However, a setup with four to six IMUs and a monocular camera can effectively achieve stable and virtually drift-free motion capture. Our experiment analysis offers substantial experimental evidence that paves the way for further research in this field. Furthermore, MINIONS can also be a benchmark dataset for many other tasks such as 2D-to-3D pose estimation [74, 47], fine-grained action recognition [18, 63], etc. The evaluation results of these tasks further exploit the potential of our dataset.
@@ II Related Work
Motion Capture from Monocular Video [31, 70, 71]. Motion capture, i.e., 3D pose estimation, from monocular videos has been a popular topic in the computer vision community [37]. Recent approaches [19, 11] can be divided into optimization-based methods and regression-based ones. Optimization-based methods [5, 21, 52] optimized parameters of the model, e.g., SMPL, to fit the joints or silhouettes of human bodies in video frames. Regression-based methods, e.g., HMR [29] and ROMP [53], learned deep neural networks from large-scale data to regress the model parameters given a single RGB image. Although these methods have achieved excellent accuracy and real-time performance on these benchmarks, they may still generate poor results and temporal jitters due to occlusion, fast actions, and subtle movements in real-world scenarios. Therefore, this paper aims to exploit sparse IMUs as supplementary cues to overcome these challenges.
@@
Motion Capture by Combination Schemes [10]. Motion capture from both videos and IMUs has also attracted the attention of researchers. Existing methods [23, 33, 51] effectively improve the motion capture accuracy by eliminating multiplayer ambiguity in videos and minimization of location drift of IMUs. For example, Gilbert et al. [20] fused multi-channel volumetric data from multi-view cameras and IMU signals to estimate 3D joints. They also built a dataset, TotalCapture [56], as listed in TABLE I. However, TotalCapture only contained 50-minute videos of five types of actions performed by five subjects, limiting the scalability of the dataset and related methods. During the data collection of TotalCapture, actors are mandated to wear tight-fitting attire, resulting in unavoidable discrepancies between the attire worn in daily life and that of the experimental setting. Therefore, we focus on building a large-scale motion capture dataset of daily life actions containing both RGB videos and IMUs and providing a more practical motion capture paradigm using monocular videos and sparse inertial signals for daily applications.
@@ IV-B Network Structure
TABLE II: Comparison between IMUs-based, monocular vision-based, and multi-modal human motion capture. #IMUs: the number and placement of IMUs used in the algorithms. In detail, 4 means {LeftHand, RightHand, Head, Hip}, 6 means {LeftForeArm, RightForeArm, LeftLeg, RightLeg, Head, Hip}, 8 means {LeftHand, RightHand, LeftForeArm, RightForeArm, LeftFoot, RightFoot, Head, Hip}, 10 means {LeftHand, RightHand, LeftForeArm, RightForeArm, LeftFoot, RightFoot, LeftLeg, RightLeg, Head, Hip}. #Cams: the number of camera views used in the algorithms. $\mu_{glo}$ and $\sigma_{glo}$: the mean and variance global rotation error of all body joints in degrees. MPJPE: the mean Euclidean distance between the predicted 3D joint positions and the corresponding ground truth joint positions. Jitter: the average jerk of body joints.
@@ V Experiments
V-A Implementation Details
V-B Multi-modal Human Motion Capture
@@
Comparison. Our experimental results are detailed in Table II. The results from our experiments show that motion capture using the monocular camera is frequently hampered by instability resulting from blurriness or occlusion and leads to large Jitter. In contrast, motion capture systems that rely on IMUs are prone to larger errors in global rotations and Euclidean distance, attributable to the inherent unreliability of the IMUs. However, an integrated system combining just four to six IMUs with a monocular camera can successfully ensure stable motion capture with minimal error and virtually no jitter. Experimental results show that a motion capture system equipped with a monocular camera and more than eight IMUs is unnecessary. They provide negligible improvements in global rotation error and Euclidean distance error, while leading to increased jitter and equipment cost. Furthermore, we conduct experiments for motion capture from a monocular camera and full-body IMUs on the TotalCapture [56] dataset in Table III. Our method outperforms other methods in terms of performance and demonstrates greater flexibility, allowing for adaptation to various sensor configurations.
@@ V-C Benchmarks on other Tasks
In this subject, we evaluate mainstream baselines and benchmarks on tasks including 2D-to-3D pose estimation [42] and Fine-grained Video Action Recognition [54] , which facilitate more comprehensive explorations of our dataset.
TABLE IV: 2D-to-3D pose estimation. #Num: Receptive Field (frames).
@@
2D-to-3D pose estimation. The 2D-to-3D pose estimation task aims to accurately predict the location of 3D joints from the detected 2D human joints, which is a specific type of 3D pose estimation task. The challenge is that depth information is lost in the projection from 3D joints to 2D joints, making the reverse process of deducing 3D joints from 2D observations ill-posed. In TABLE IV, we evaluate the accuracy of the mainstream methods. MPJPE is applied for measuring the average Euclidean distance between the ground truth and predicted joint positions, while PA-MPJPE is a similar metric that first aligns the predicted pose to the ground truth using a rigid transformation before computing the error. To explore the upper bounds of current mainstream methods, we use the groundtruth of 2D joints as input. The experimental results indicate that there is still room for improvement in this task.
Fine-grained Video Action Recognition. Our dataset provides fine-grained action labels and motion videos from various views, which can be validated for applications in fine-grained video action recognition. As shown in TABLE V, we test the accuracy of the mainstream video action understanding methods on our dataset. Following these methods, Top-1 and Top-5 are used to evaluate model performance. Top-1 is the model’s accuracy in predicting the most likely action, while Top-5 is the accuracy at which the true action is within the model’s top-five predictions. Compared to the mainstream video understanding datasets, e.g. Kinetics400 [7], NTU RGB+D 60 [50], and NTU RGB+D 120 [35], our dataset is more challenging since existing methods easily achieve Top-1 accuracy of more than 85% on the Kinetics400 dataset.
TABLE V: Fine-grained Video Action Recognition on datasets.
@@
VI Conclusion

### v4

@@ I Introduction
Human motion capture is the process of recording human movement represented by a sequence of 3D positions and rotations of mesh or joints of the human body [29, 12, 11]. Industrial motion capture systems have been widely applied in movie and game production, sports analysis, medical diagnosis, etc. However, these systems usually consist of tens of synchronized cameras or a group of wearable sensors and specific signal receivers [1, 2, 68]. Despite their high accuracy for human motion capture, individual consumers can hardly afford the high cost and learn professional configurations. Therefore, this paper investigates accurate multi-modal human motion capture with consumer-affordable devices and easy-to-use operations for daily applications [85, 46, 13, 43] like eXtended Reality (XR), live video streaming, etc.
@@
Based on MINIONS, we propose a two-branch framework, named SparseNet, for motion capture by fusion of videos and IMUs. The core of SparseNet is a deliberately designed two-branch structure grounded in Bayesian probability theory and Matrix Fisher priors [15, 32]. For the sparse IMU signals, the Sparse IMUs branch takes inertial measurements as input to learn posture features, providing rotational probability distribution. To overcome the incompleteness and location drift of sparse IMUs, we design a visual branch to supplement the missed joints and correct the location drift, estimating 3D bone probability distribution and pose probability prior. This framework enables adaptive integration through a posterior fusion module that aligns distributions from both branches. Experiments reveal that the visual branch for monocular motion capture is often limited by jittering due to blurriness or occlusion, while the IMUs branch tends to suffer from global position drifts. However, a setup with four to six IMUs and a monocular camera can effectively achieve stable and virtually drift-free motion capture. Our experiment analysis offers substantial experimental evidence that paves the way for further research in this field. Furthermore, MINIONS can also be a benchmark dataset for many other tasks such as 2D-to-3D pose estimation [87, 57], fine-grained action recognition [20, 76], gait activity assessment, etc. The evaluation results of these tasks further exploit the potential of our dataset.
@@ II Related Work
Motion Capture from Monocular Video [33, 83, 42]. Motion capture, i.e., 3D pose estimation, from monocular videos has been a popular topic in the computer vision community [44]. Recent approaches [21, 40, 39] can be divided into optimization-based methods and regression-based ones. Optimization-based methods [6, 24, 64] optimized parameters of the model, e.g., SMPL, to fit the joints or silhouettes of human bodies in video frames. Regression-based methods, e.g., HMR [31] and ROMP [65], learned deep neural networks from large-scale data to regress the model parameters given a single RGB image. Although these methods have achieved excellent accuracy and real-time performance on these benchmarks, they may still generate poor results and temporal jitters due to occlusion, fast actions, and subtle movements in real-world scenarios. Therefore, this paper aims to exploit sparse IMUs as supplementary cues to overcome these challenges.
@@
Motion Capture by Combination Schemes. Motion capture from both videos and IMUs has also attracted the attention of researchers. Existing methods [26, 36, 63] effectively improve the motion capture accuracy by eliminating multiplayer ambiguity in videos and minimization of location drift of IMUs. For example, Gilbert et al. [22] fused multi-channel volumetric data from multi-view cameras and IMU signals to estimate 3D joints. They also built a dataset, TotalCapture [68], as listed in TABLE I. However, TotalCapture only contained 50-minute videos of five types of actions performed by five subjects, limiting the scalability of the dataset and related methods. During the data collection of TotalCapture, actors are mandated to wear tight-fitting attire, resulting in unavoidable discrepancies between the attire worn in daily life and that of the experimental setting. Therefore, we focus on building a large-scale motion capture dataset of daily life actions containing both RGB videos and IMUs and providing a more practical motion capture paradigm using monocular videos and sparse inertial signals for daily applications.
@@ IV-B Network Structure
TABLE II: Comparison between IMUs-based, monocular vision-based, and multi-modal human motion capture. #IMUs: the number and placement of IMUs used in the algorithms. In detail, 2 means {Head, Hip}, 4 means {LeftHand, RightHand, Head, Hip}, 6 means {LeftForeArm, RightForeArm, LeftLeg, RightLeg, Head, Hip}, 8 means {LeftHand, RightHand, LeftForeArm, RightForeArm, LeftFoot, RightFoot, Head, Hip}, 10 means {LeftHand, RightHand, LeftForeArm, RightForeArm, LeftFoot, RightFoot, LeftLeg, RightLeg, Head, Hip}. #Cams: the number of camera views used in the algorithms. $\mu_{glo}$ and $\sigma_{glo}$: the mean and variance global rotation error of all body joints in degrees. MPJPE: the mean Euclidean distance between the predicted 3D joint positions and the corresponding ground truth joint positions. Jitter: the average jerk of body joints.
@@ V Experiments
V-A Implementation Details
V-B Multi-modal Human Motion Capture
TABLE IV: FPS comparison across sensing paradigms. For IMU-based and multi-modal settings, we report FPS under 2/4/6/8/10 IMUs, respectively. “#Cams” denotes the number of cameras.
Paradigm #IMUs #Cams FPS $\uparrow$
IMU-based 2 0 522
4 0 515
6 0 507
8 0 501
10 0 492
Vision-based 0 1 87
Multi-modal 2 1 76
4 1 75
6 1 75
8 1 74
10 1 74
Comparison. Our experimental results are detailed in Table II. The results from our experiments show that motion capture using the monocular camera is frequently hampered by instability resulting from blurriness or occlusion and leads to large Jitter. In contrast, motion capture systems that rely on IMUs are prone to larger errors in global rotations and Euclidean distance, attributable to the inherent unreliability of the IMUs. However, an integrated system combining just four to six IMUs with a monocular camera can successfully ensure stable motion capture with minimal error and virtually no jitter. We further evaluate an ultra-sparse setting with only two IMUs ({Head, Hip}), which shows a clear performance degradation due to insufficient inertial constraints, but our multi-modal fusion still produces relatively stable motion and suppresses jitter compared to the IMU-only baseline. In addition, we report the inference speed (FPS) in Table IV, showing that our approach runs comfortably above real-time. Experimental results show that a motion capture system equipped with a monocular camera and more than eight IMUs is unnecessary. They provide negligible improvements in global rotation error and Euclidean distance error, while leading to increased jitter and equipment cost. Furthermore, we conduct experiments for motion capture from a monocular camera and full-body IMUs on the TotalCapture [68] dataset in Table III. Our method outperforms other methods in terms of performance and demonstrates greater flexibility, allowing for adaptation to various sensor configurations.
Computational Cost and Runtime. We report the end-to-end inference speed (FPS) for different sensing paradigms. All measurements are conducted on a single RTX 4090 GPU. As shown in Table IV, our method runs at 492–522 FPS in the IMU-based setting and 74–76 FPS in the multi-modal setting (IMU+1 camera), which is well above real-time requirements (e.g., 30 FPS). The computational cost is largely independent of action dynamics (slow vs. highly dynamic) and scales approximately proportionally with the number of tracked persons.
@@ V-C Benchmarks on other Tasks
In this subject, we evaluate mainstream baselines and benchmarks on tasks including 2D-to-3D pose estimation [49], Fine-grained Video Action Recognition [66], In-the-wild Monocular Pose Estimation Evaluation [47] and Gait Activity Assessment, which facilitate more comprehensive explorations of our dataset.
TABLE V: 2D-to-3D pose estimation. #Num: Receptive Field (frames).
@@
2D-to-3D pose estimation. The 2D-to-3D pose estimation task aims to accurately predict the location of 3D joints from the detected 2D human joints, which is a specific type of 3D pose estimation task. The challenge is that depth information is lost in the projection from 3D joints to 2D joints, making the reverse process of deducing 3D joints from 2D observations ill-posed. In TABLE V, we evaluate the accuracy of the mainstream methods. MPJPE is applied for measuring the average Euclidean distance between the ground truth and predicted joint positions, while PA-MPJPE is a similar metric that first aligns the predicted pose to the ground truth using a rigid transformation before computing the error. To explore the upper bounds of current mainstream methods, we use the groundtruth of 2D joints as input. The experimental results indicate that there is still room for improvement in this task.
Fine-grained Video Action Recognition. Our dataset provides fine-grained action labels and motion videos from various views, which can be validated for applications in fine-grained video action recognition. As shown in TABLE VI, we test the accuracy of the mainstream video action understanding methods on our dataset. Following these methods, Top-1 and Top-5 are used to evaluate model performance. Top-1 is the model’s accuracy in predicting the most likely action, while Top-5 is the accuracy at which the true action is within the model’s top-five predictions. Compared to the mainstream video understanding datasets, e.g. Kinetics400 [8], NTU RGB+D 60 [60], and NTU RGB+D 120 [38], our dataset is more challenging since existing methods easily achieve Top-1 accuracy of more than 85% on the Kinetics400 dataset.
TABLE VI: Fine-grained Video Action Recognition on datasets.
@@
In-the-wild Monocular Pose Estimation Evaluation. We additionally evaluate our approach on the in-the-wild [52] monocular benchmark 3DPW [47]. We report the commonly used metrics (all in mm): Procrustes-aligned MPJPE (PA-MPJPE) and MPJPE. As shown in Table VII, our method achieves 35.4 PA-MPJPE and 56.6 MPJPE, indicating good generalization beyond the controlled capture setup.
TABLE VII: In-the-wild monocular performance comparison on 3DPW [47]. We report the commonly used metrics (all in mm): Procrustes-aligned MPJPE (PA-MPJPE) and MPJPE.
Method PA-$\downarrow MPJPE\downarrow$
BEDLAM-CLIFF [5] 46.6 72.0
CLIFF [35] 43.0 69.0
HMR2.0 [23] 44.4 69.8
TokenHMR [17] 44.3 71.0
PromptHMR [72] 36.6 58.7
WHAM [62] 37.5 59.8
GVHMR [61] 37.0 56.6
TRAM [73] 35.6 59.3
Ours 35.4 56.6
Part-wise Lower-limb Analysis for Gait Activity Assessment. To better understand the gait-related potential [54, 53] of MINIONS, we provide a part-wise analysis on lower-limb components (feet/legs/upper-legs). We report average results over joints belonging to LEFTFOOT, RIGHTFOOT, LEFTLEG, RIGHTLEG, LEFTUPLEG, and RIGHTUPLEG. As shown in Table VIII, our multi-modal model consistently improves lower-limb pose accuracy over the IMU-only baseline across all parts.
TABLE VIII: Part-wise lower-limb analysis on MINIONS for gait activity assessment. We report global orientation error statistics ($\mu_{glo}$$\sigma_{glo}$, deg), MPJPE ($m$), and Jitter ($10^{2}m/s^{3}$), where lower is better. We show results of an IMU-only baseline (IMUs-based) and our full multi-modal model.
[…]
IMUs-based LEFTFOOT 13.84 10.26 62.57 1.45
RIGHTFOOT 13.63 10.07 61.86 1.42
LEFTLEG 10.93 8.04 49.57 1.23
RIGHTLEG 10.74 7.96 48.91 1.18
LEFTUPLEG 9.67 7.13 44.25 1.05
RIGHTUPLEG 9.58 7.09 43.82 1.03
Multi-modal LEFTFOOT 10.27 6.86 46.07 1.85
RIGHTFOOT 10.08 6.67 45.43 1.82
LEFTLEG 8.47 5.78 37.25 1.64
RIGHTLEG 8.39 5.68 36.82 1.63
LEFTUPLEG 7.64 5.16 33.57 1.48
RIGHTUPLEG 7.59 5.07 33.11 1.46
VI Ethical Considerations and Dataset Access
The principal ethical consideration addressed in this work is privacy, since the dataset contains multi-modal sensor signals and RGB videos that may include personally identifiable appearance information. We take privacy protection seriously and are committed to safeguarding the confidentiality of participants’ data.
Informed consent. Informed consent was obtained from all individual participants included in the study. For transparency, the full informed consent form text is provided in the supplementary material.
Controlled access. MINIONS is not intended for indiscriminate public download. The dataset is distributed via a registration-based and approval-based controlled access mechanism. Eligible users from institutions submit the required information and a signed agreement, and then receive download instructions under a Data License Agreement that restricts redistribution and requires responsible use.
Project page. The dataset access instructions are available at: https://huggingface.co/datasets/Sheldong/MINIONS.
VII Conclusion